OpenStudy (anonymous):

The x-intercepts of the graph of ƒ(x) = x(x + 1)(x – 5)(x – 7) are: (there is four answers) 1. 0 or 1 2. 1 or -1 3. 5 or -5 4. 7 or -7

OpenStudy (danjs):

x intercept is when y=0 right?

OpenStudy (danjs):

If you have terms multiplied together like: A * B * C = 0 Only one of the terms has to be zero for the equation to be true.

OpenStudy (danjs):

so in your case x=0 (x + 1) = 0 (x - 5) = 0 (x - 7) = 0

OpenStudy (danjs):

See that?

OpenStudy (anonymous):

ooooooo now i do!